If 10 20 molecules are removed from 220 mg of CO 2 then the number of moles of CO 2 left are?

If 1020  molecules are removed from 220 mg of COthen the number of moles of COleft are?

Grade:10

1 Answers

Ritik
6 years ago
It can be done by change 10^20 molecules in to mole (devide it by Avogadro Number).Mole of 10^20 molecules =0.000166Now mole of 220mg CO2= 220*10^-3/44 (molecular weight of CO2=44g)=0.005So now number of left mole= 0.005-0.000166=0.004834